A tank can be emptied thru either of 2 pipes, the 1st pipe empties the tank in 4hours. The 2nd pipe empties in 3hours. how long will it take to emptythe tank using both pipes
Found 2 solutions by stanbon, Lacey020991:
Answer by stanbon(75887)

A tank can be emptied thru either of 2 pipes, the 1st pipe empties the tank in 4hours. The 2nd pipe empties in 3hours. how long will it take to emptythe tank using both pipes
-----------------------
1st Pipe DATA:
Time = 4 hours/job ; Rate = 1/4 job/hr
--------------
2nd Pipe DATA:
Time = 3 hours/job ; Rate = 1/3 job/hr
--------------
Together DATA;
Time = x hours/job ; Rate = 1/x job/hr
----------------
EQUATION:
rate + rate = together rate
1/4 + 1/3 =1/x
7/12 = 1/x
x= 12/7 = 1 hr (5/7)*60 minutes
x = 1 hr 42.86 minutes ( Time to empty the tank if both pipes are open)
===============
Cheers,
Stan H.

Answer by Lacey020991(49)